OpenStudy (anonymous):

what is wrong with the equation? x^2+x-6/x-2= x+3

OpenStudy (anonymous):

\[\frac{x^2+x-6}{x-2}=\frac{(x+3)(x-2)}{x-2}=x+3\] only if \(x\neq 2\)

OpenStudy (anonymous):

if \(x=2\) then \(x+3=2+3=5\) but \(\frac{x^2+x-6}{x-2}\) will be undefined since you cannot divide by 0

OpenStudy (anonymous):

makes sense! thanks! tough to learn with a Russian prof ha